### v3

# The causal effects of modified treatment policies under network interference

Abstract
Modified treatment policies are a widely applicable class of interventions useful for studying the causal effects of continuous exposures. Approaches to evaluating their causal effects assume no interference, meaning that such effects cannot be learned from data in settings where the exposure of one unit affects the outcomes of others, as is common in spatial or network data. We introduce a new class of intervention—induced modified treatment policies—which we show identify such causal effects in the presence of network interference. Building on recent developments for causal inference in networks, we provide flexible, semi-parametric efficient estimators of the statistical estimand. Numerical experiments demonstrate that an induced modified treatment policy can eliminate the causal, or identification, bias that results from network interference. We use the methodology developed to evaluate the effect of zero-emission vehicle uptake on air pollution in California, strengthening prior evidence.
1 Introduction
Scientists frequently seek to understand the causal effect of a given policy on an outcome $Y$ in a target population. Such policies commonly involve modifying the natural value of some continuous treatment or exposure $A$; that is, the value $A$ would have taken had its distribution been left unaltered (Sarvet & Stensrud, 2025). Often, the scientist’s ideal goal is to answer the question, “how would the population mean of $Y$ change if the natural value of $A$ were increased or decreased?”
Modified treatment policies (MTPs) (Robins et al., 2004; Díaz & van der Laan, 2012; Haneuse & Rotnitzky, 2013; Young et al., 2014) are a class of interventions that define causal estimands well-suited for formulating such counterfactual questions about continuous exposures. An MTP can answer the question, “how much would the average value of $Y$ have changed had the natural value of $A$ been increased or decreased by an increment $\delta$?” where $\delta$ is chosen by the investigator. Hence, the MTP provides a causally interpretable generalization of the commonly applied procedure of estimating a regression coefficient. In fact, under specific structural assumptions, the effect of applying the MTP $A+1$ to every study unit is equivalent to the coefficient for the exposure $A$ in a main terms linear regression. Hence, though rarely mentioned explicitly by name, the causal effect of an MTP is a popular target estimand in studies that aim to evaluate policy effects on a well-defined population.
The causal effect of an MTP carries several key advantages over alternative estimands. Unlike the causal dose-response curve, the MTP effect is identified even in settings where it may not be feasible—or, indeed, sensible—to deterministically set each unit’s exposure to the same level. Consequently, the MTP effect is identified even when certain exposure levels may suffer from non-overlap among sub-populations. Furthermore, the MTP effect estimand admits standard semi-parametric efficient estimators, facilitating the use of machine learning to flexibly capture nonlinear relationships and thus allowing the investigator to mitigate the risks posed by model misspecification.
Like other common causal inference methods, a key assumption of the classical MTP framework is non-interference—that one unit’s exposure does not impact any other units’ outcome (Cox, 1958; Rubin, 1980). This critical assumption cannot be made when, for example, study units correspond to geographic areas, such as counties or ZIP code areas. Here, the population units may move around; thus, assigning an exposure policy to a given county would affect not only those residing in that geographic region but also those commuting to and from said region. In general, ignoring interference between units invalidates identification and overtly introduces bias into analytic results (Halloran & Hudgens, 2016). Despite these challenges, data from studies that involve interference between study units can be exceedingly useful, even critical, to advancing science and policy in many settings of current interest. For example, environmental epidemiologists routinely leverage observational spatial data to study the effects of pollution on human health (Elliott & Wartenberg, 2004; Reich et al., 2021; Morrison et al., 2024), a setting in which randomization is impractical and unethical, and where the “shifts in distribution” measured by MTPs are of interest (Tec et al., 2024).
Owing to its straightforward interpretability and the non-restrictive conditions for its identification, especially when compared to other causal estimands for continuous exposures, it is useful to deepen understanding of how to apply the MTP framework even in settings when non-interference cannot be assumed. The present work develops a framework for both identifying and efficiently estimating an MTP’s causal effect under network interference. This allows investigators to obtain inference about the causal effect of a continuous exposure while both accounting for network interference between units and incorporating flexible regression procedures in the estimation process.
Contributions. We present several contributions to the literature on causal inference under network interference. Firstly, we introduce the concept of an induced MTP, a new type of intervention that identifies the causal effect of an MTP when network interference is present. This intervention accounts for how the application of an MTP to a given unit affects its neighbors in the known (or assumed) network structure. Secondly, applying the “coarea formula” from measure-theoretic calculus, we provide a novel identification result for the causal effect of an induced MTP in the network setting. Finally, we develop semi-parametric efficient estimators for the causal effect of an induced MTP by applying and building on the prior theoretical contributions of Ogburn et al. (2022). Specifically, we resolve several of their methodological challenges, including (1) using cross-fitted machine learning for nuisance parameter estimation in place of restrictive parametric regression strategies; (2) deriving more tractable forms of the relevant nuisance parameters, which can be reliably estimated using standard regression tools; (3) eliminating reliance on computationally-intensive Monte Carlo procedures for estimation and inference; and (4) obtaining consistent variance estimators.
Outline. Section 2 reviews MTPs and causal inference under interference. Section 3 describes the induced MTP, including identification and estimation of causal effects. Section 4 reports numerical results verifying the proposed methodology, while Section 5 discusses an illustrative data analysis in our motivating applied science context. We conclude and discuss future directions in Section 6.
2 Background
Throughout, we let capital bold letters denote random $n$-vectors; for instance, $\mathbf{Y}=(Y_{1},\ldots,Y_{n})$. Consider data $\mathbf{O}=(\mathbf{L},\mathbf{A},\mathbf{Y})\sim\mathsf{P}\in\mathcal{M}$, where $\mathsf{P}$ is the true and unknown data-generating distribution of $\mathbf{O}$ and $\mathcal{M}$ is a non-parametric statistical model (that is, a set of candidate data-generating probability distributions) that places no restrictions on the data-generating distribution. In principle, $\mathcal{M}$ may be restricted to incorporate any available real-world knowledge about the system under study. Let $O_{i}=(L_{i},A_{i},Y_{i})$ represent measurements on the ith individual data unit, where $Y_{i}$ is an outcome of interest, $A_{i}$ is a continuous exposure with support $\mathcal{A}$, and $L_{i}$ is a collection of baseline (that is, pre-exposure) covariates. To ease notational burden, we omit subscripts $i$ when referring to an arbitrary unit $i$ (that is, $Y\equiv Y_{i}$ when the specific unit index is taken to be uninformative).
We will assume that the data-generating process can be expressed via a structural causal model (SCM, Pearl (2000)) encoding the temporal ordering between variables: $\mathbf{L}$ is generated first, then $\mathbf{A}$, and finally $\mathbf{Y}$. We denote by $Y(a)$ the counterfactual random variable (or potential outcome) generated by hypothetically intervening upon $A$ to set it to $a\in\mathcal{A}$ and allowing the impact of such an intervention to propagate downstream, to the component of the SCM that generates $Y$. Our goal is to reason about the causal relationship between $\mathbf{A}$ and $\mathbf{Y}$ in spite of the presence of confounders $\mathbf{L}$ and network interference between units $i=1,\ldots,n$.
2.1 Continuous Exposures with Modified Treatment Policies
A Modified Treatment Policy (MTP) is a user-specified function $d(a,l;\delta)$ that maps the observed value $a$ of an exposure $A$ to a new value and may itself depend on the natural (or pre-intervention) value of the exposure (Haneuse & Rotnitzky, 2013).
Example 1 (Additive Shift).
For a fixed $\delta$, an additive shift MTP may be defined as
$d(a,l;\delta)=a+\delta\ .$ (1)
This corresponds to the scientific question, “how much of a change in $Y$ would be caused by adding $\delta$ to the observed natural value of $a$, for all units regardless of their stratum ll?”
Example 2 (Multiplicative Shift).
For a fixed $\delta$, a multiplicative shift MTP is defined as
$d(a,l;\delta)=\delta\cdot a\ .$ (2)
This asks the scientific question, “how much of a change in $Y$ would be caused by scaling the observed natural value of $a$ by $\delta$, for all units regardless of their stratum ll?”
Note that in the above, $\delta$ is a fixed, user-specified parameter specifying the magnitude of the hypothetical intervention. However, the MTP framework is even more flexible: interventions that change depending on the values of measured covariates $L$ are allowed.
Example 3 (Piecewise Additive Shift).
Consider a piecewise additive function
$d(a,l;\delta)=\begin{cases}a+\delta\cdot l&a\in\mathcal{A}(l)\\ a&\text{otherwise}\ ,\end{cases}$ (3)
which applies an intervention whose scale depends on the value of a covariate $l$ and only occurs if the natural exposure value $a$ is within some specific subset $\mathcal{A}(l)\subset\mathcal{A}$ of the support of $A$.
MTPs can be used to define scientifically relevant causal estimands for continuous exposures. The population intervention causal effect of an MTP (Díaz & van der Laan, 2012) is defined as $\mathbb{E}_{\mathsf{P}}[Y(d(A,L;\delta))-Y]$; that is, the average difference between the outcome $Y$ that did occur under the observed natural value of treatment $A$ and the counterfactual outcome $Y(d(A,L;\delta))$ that would have occurred under the investigator-supplied MTP $d(A,L;\delta)$. This estimand answers the scientific question, “what would happen if we applied, to the study population, a policy that modified the existing exposure according to a rule encoded by $d(\cdot;\delta)$?”
When $\delta=1$, the additive MTP (Example 1) carries the familiar interpretation often attributed to a linear regression coefficient; when $\delta=1.01$, the multiplicative MTP (Example 2) holds the same interpretation, but for a log-transformed exposure. Hence, MTPs may be seen as a non-parametric and causally interpretable extension of widely used associational estimands, formalizing the problem of quantifying how the mean counterfactual outcome would change under a shift in exposure value. MTPs allow the investigator to specify a wide range of interpretable interventions on continuous exposures that may be carried out in practice. The MTP framework has gained traction for its applicability in settings involving longitudinal, time-varying interventions (Díaz et al., 2023; Hoffman et al., 2024); causal mediation analysis (Díaz & Hejazi, 2020; Hejazi et al., 2023), including with time-varying mediator–confounder feedback (Gilbert et al., 2024); and causal survival analysis under competing risks (Díaz et al., 2024). However, no work has, to our knowledge, extended the MTP framework to settings with dependent data characterized by interference between units.
2.2 Causal Inference Under Interference
Interference occurs when, for a given unit $i$, the outcome of interest $Y_{i}$ depends not only on its own assigned exposure $A_{i}$ but also upon the exposure $A_{j}$ of at least one other unit ($j\neq i$). Formally, we say interference occurs when $Y(a_{i},a_{j})\neq Y(a_{i},a_{j}^{\prime})$ if $a_{j}\neq a_{j}^{\prime}$. It is a component of the well-known stable unit treatment value assumption (SUTVA; Rubin, 1980), commonly assumed for identification of causal estimands, including those of MTPs (Haneuse & Rotnitzky, 2013; Young et al., 2014).
Previous work has focused on settings exhibiting partial interference (Hudgens & Halloran, 2008; Tchetgen Tchetgen & VanderWeele, 2012; Halloran & Hudgens, 2016), which occurs when units can be partitioned into clusters such that interference only occurs between units in the same cluster. Our work focuses instead on a broader setting, that of network interference (van der Laan, 2014), which occurs when a unit’s outcome is subject to interference by other units’ exposures according to some arbitrary known network of relationships between units. When such interference is present, the data $\mathbf{O}$ includes an adjacency matrix or network profile, $\mathbf{F}$, describing each unit’s neighboring units, occasionally termed “friends” (Sofrygin & van der Laan, 2017).
Aronow & Samii (2017) demonstrate how to identify a causal estimand when SUTVA is violated due to network interference. To do this, they use an exposure mapping: a function that maps the exposure assignment vector $\mathbf{A}$ to the exposure actually received by each unit. The exposure received is a function of a unit’s original exposure $\mathbf{A}$ and covariates $\mathbf{L}$, including the network profile $\mathbf{F}$. If the exposure mapping is correctly specified and consistent, then SUTVA is restored, and causal effects subject to interference can be identified for the exposure arising under the exposure mapping.
Ogburn et al. (2022) and van der Laan (2014) rely on similar logic to identify population causal effects from data exhibiting a causally dependent structure, doing so by constructing “summary functions” of neighboring units’ exposures. Notably, they describe semi-parametric theory for efficiently estimating the effects of stochastic interventions in the network dependence setting. Stochastic interventions, which differ from MTPs, replace the natural value of exposure with a random draw from an investigator-supplied counterfactual distribution (Díaz & van der Laan, 2012). While a mathematically elegant strategy, the interpretation of a stochastic intervention is typically challenging—at times, even impractical—as real-world policies can seldom be defined by randomly assigning (post-intervention) exposure values to study units. Furthermore, this setting gives rise to estimation challenges: empirical evaluations in Ogburn et al. (2022) and related work by Sofrygin & van der Laan (2017) relied on relatively bespoke parametric modeling of nuisance parameters, coupled with Monte Carlo procedures for point and variance estimation.
Notably, while the hypothetical exposure that results from this random draw is not guaranteed to match that which would result from an MTP, the two classes of interventions may be constructed to yield equivalent counterfactual means (Young et al., 2014). Given the similarities between these intervention schemes, we extend recent theoretical developments to construct semi-parametric efficient estimators of the effects of MTPs under network interference. Our work reveals that, for MTPs, much of the previously established semi-parametric theory can be reduced in ways that simplify the application of machine learning and adaptive, non-parametric regression for nuisance estimation.
But what does it mean to intervene on a summary function? If one were to intervene on the summary directly, the resulting collection of counterfactual exposures could plausibly be inconsistent with the structure of the network. Some existing works seek to circumvent this issue by recasting the desired estimand as a mean of individual-level causal effects (Aronow & Samii, 2017; Athey et al., 2017; Sävje, 2023). When investigators aim to estimate the impact of a hypothetical policy, however, this strategy will not answer questions of scientific interest—for the estimand does not correspond to a population-level intervention that could be implemented in practice. In order to estimate a population-level causal effect, one must consider first intervening, and only then applying the summary or exposure mapping—a process more naturally applicable to MTPs.
Other relevant works address interference under different assumptions: random networks (Clark & Handcock, 2024), multiple outcomes (Shin et al., 2023), long-range dependence (Tchetgen Tchetgen et al., 2021; Liu et al., 2025), bipartite graphs (Zigler et al., 2023), and unknown network structure (Ohnishi et al., 2022; Hoshino & Yanagi, 2023). We build on the setting described by Ogburn et al. (2022), as their scientific goals most closely resemble those of the MTP framework.
3 Methodology
Suppose there exists a network describing whether two units are causally dependent with adjacency matrix $\mathbf{F}$, where $F_{i}$ denotes the neighbors of unit $i$. For each unit $i$, a set of confounders $L_{i}$ is drawn, followed by an exposure $A_{i}$ based on a summary $L_{i}^{s}$ of its own and its neighbors’ confounders, and finally an outcome $Y_{i}$ based on $L_{i}^{s}$ and a summary of its own and its neighbors’ exposures, $A_{i}^{s}$. This data-generating process can be defined formally as the SCM in Equation (4):
$L_{i}=f_{L}(\varepsilon_{L_{i}});A_{i}=f_{A}(L_{i}^{s},\varepsilon_{A_{i}});Y_{i}=f_{Y}(A_{i}^{s},L_{i}^{s},\varepsilon_{Y_{i}})\ .$ (4)
Since $A_{i}^{s}$ and $L_{i}^{s}$ are not directly observed but are functions that follow from the scientific problem at hand, it follows that (in a slight abuse of notation) $A_{i}^{s}=s_{F_{i}}(\mathbf{A},\mathbf{L})$ and $L_{i}^{s}=s_{F_{i}}(\mathbf{L})$, which we use to denote that, for unit $i$, the summary function depends on $i$’s column of the network adjacency matrix $\mathbf{F}$. This means that $A_{i}^{s}$ and $L_{i}^{s}$ depend only on neighboring units. As just one example, $A_{i}^{s}$ may be a (possibly weighted) sum of neighbors: $s_{F_{i}}(\mathbf{A},\mathbf{L})=\sum_{j=1}^{n}F_{ij}A_{j}=\sum_{F_{ij}\neq 0}A_{j}$.
Following Ogburn et al. (2022), we assume error vectors $(\varepsilon_{L_{1}},\ldots,\varepsilon_{L_{n}})$, $(\varepsilon_{A_{1}},\ldots,\varepsilon_{A_{n}})$, and $(\varepsilon_{Y_{1}},\ldots,\varepsilon_{Y_{n}})$ are independent of each other, with entries identically distributed and either $\varepsilon_{i}\mbox{$\perp\!\!\!\perp$}\varepsilon_{j}$ provided $\{i,j\}\not\subseteq F_{k},\,\forall\,\,k\in 1,\ldots,n$ or $\text{Cov}(\varepsilon_{i},\varepsilon_{j})\geq 0$ otherwise. That is, errors between units are independent provided that the units are neither directly connected nor share ties with a common node in the interference network given by $\mathbf{F}$; otherwise, errors across units may be positively correlated. Positive correlation ensures Theorem 1 will hold, and is typical in applied settings with interference.
Interference bias arises when the data arise from the SCM (4) but investigators wrongly assume that $f_{Y}$ is a function only of $A_{i}$ and $L_{i}$, and not of $\{A_{j}\colon F_{ij}\neq 0\}$ or $\{L_{j}\colon F_{ij}\neq 0\}$. Since interference violates the consistency rule (Pearl, 2010), commonly relied upon to identify causal effects, ignoring its presence, even inadvertently, leads to a failure in identification and consequently risks biased estimation. Under the SCM (4), identifiability of the causal effect of applying an exposure to all units $A_{j}\colon j\in F_{i}$ can be restored by controlling for all $L_{j}\colon F_{ij}\neq 0$ directly or via the dimension-reducing summaries $A_{i}^{s}$ and $L_{i}^{s}$ (van der Laan, 2014) of a unit’s neighbors’ confounders and exposures.
3.1 Induced Modified Treatment Policies
As discussed in Section 2, restoring identifiability by performing inference on $A^{s}$ and $L^{s}$ instead of $A$ and $L$ is a common strategy for handling interference. This approach is, however, not compatible with the application of MTPs: one must consider not the causal effect of $A$ under $d(\cdot;\delta)$; but rather, the effect of $A^{s}$ after intervening on the upstream exposure via $d(\cdot;\delta)$. To identify the causal effects of MTPs under interference, we introduce a novel intervention scheme—the induced MTP.
Consider applying an MTP to the SCM (4), replacing $\mathbf{A}$ with the vector $\mathbf{A}^{d}=d(\mathbf{A},\mathbf{L};\delta)=[d(A_{i},L_{i};\delta)]_{i=1}^{n}$. Under interference, we are interested in the causal effect of $A^{s}$ on $Y$. Hence, the scientific question of interest is actually, “what if $A_{i}^{s}$ were replaced by some $A_{i}^{s\circ d}=s_{F_{i}}(d(\mathbf{A},\mathbf{L};\delta),\mathbf{L})$?” This process is illustrated in Figure 1. We call the function composition $s\circ d$ the induced MTP.
$\mathbf{A}$$\mathbf{A}^{d}$$\mathbf{A}^{s\circ d}$ddss
Figure 1: How an induced MTP arises as the composition of MTP and summary functions $d\circ s_{A}$.
The counterfactual mean of an induced MTP is given by Equation (5):
$\Psi_{n}(\mathsf{P})=\mathbb{E}_{\mathsf{P}}\Big[\frac{1}{n}\sum_{i=1}^{n}Y(s_{F_{i}}(d(\mathbf{A},\mathbf{L};\delta),\mathbf{L}))\Big]\ .$ (5)
Under an induced MTP, interference no longer hampers identifiability because $s_{F_{i}}(\mathbf{A},\mathbf{L})$ captures the contribution of all relevant units (that is, a given unit $i$ and its neighbors) to each $Y_{i}$. This data-adaptive parameter will converge to the population counterfactual mean as $n\rightarrow\infty$ (Hubbard et al., 2016). Use of such a parameter definition is necessary because we must condition on the single observation of the interference network at play. Hence, do note that in all theory that follows, our estimates will implicitly condition on the observed network $\mathbf{F}$. This estimand must be compatible with the network: it is interpreted as the average change in $Y$ caused by imposing the unit-level MTP $d$ on each unit in the population governed by the network $\mathbf{F}$. With $\Psi_{n}(\mathsf{P})$ identified, the population intervention effect (a contrast) may be defined by subtracting $\mathbb{E}Y$ (Díaz & van der Laan, 2012).
3.2 Identification
In addition to the SCM (4), to identify the causal parameter $\Psi_{n}(\mathsf{P})$ by a statistical parameter $\psi_{n}$, we rely on the following assumptions:
A1Summary positivity.
If the summaries satisfy $(s_{F_{i}}(\mathbf{a},\mathbf{l}),s_{F_{i}}(\mathbf{l}))\in\text{supp}\{A_{i}^{s},L_{i}^{s}\}$, then the summaries after intervention satisfy $(s_{F_{i}}(\mathbf{a}^{d},\mathbf{l}),s_{F_{i}}(\mathbf{l}))\in\text{supp}\{A_{i}^{s},L_{i}^{s}\}$.
Assumption A1 is a weaker positivity requirement than that required for identification of the causal dose-response curve—rather than requiring all possible exposure values to be observable for every combination of covariates, we only require that the MTP keep the exposure of each unit within the support defined by its own covariates, that is, for its own stratum. To some extent, this can be enforced by design if the investigator were to choose $d$ appropriately. Furthermore, Assumption A1 is even weaker than the standard MTP positivity assumption: technically, $A_{j}^{d}\in\text{supp}\{A,L\}$ for all $j\in F_{i}$ is not required, only that the summary $A_{i}^{s\circ d}$ remains in the support of $A_{i}^{s}$. In other words, we do not require positivity for individual exposures, only positivity on the summary as a whole. Importantly, $\text{supp}\{A_{i}^{s},L_{i}^{s}\}$ denotes the support of $A_{i}^{s}$ and $L_{i}^{s}$ together implied by the number of neighbors $F_{i}$, as, under a fixed network adjacency matrix $\mathbf{F}$, the summary function that produces $A_{i}^{s}$ may depend on $\mathbf{L}$ and $F_{i}$. Since $\text{supp}\{A_{i}^{s},L_{i}^{s}\}$ may differ for each unit depending on the number of neighbors, this could be much smaller than $\mathcal{A}_{i}^{s}\bigotimes\mathcal{L}_{i}^{s}$ for certain summaries.
A2No unmeasured confounding.
The potential outcome satisfies the condition $Y(s_{F_{i}}(\mathbf{a},\mathbf{L}))\mbox{$\perp\!\!\!\perp$}s_{F_{i}}(\mathbf{A},\mathbf{L})\mid\mathbf{L}$ for all $\mathbf{a}\in(\mathcal{A})^{n}$ and $i\in\{1,\ldots,n\}$ .
This assumption may be interpreted as slightly weaker than the typical no unmeasured confounding assumption in causal inference: we only require that potential outcomes of a unit $i$ are independent of all possible exposure summaries that could be obtained conditional on the neighbors involved in the summary. For example, if a particular unit’s summary only depends on a single neighbor’s exposure, then for that unit, the analysis need only adjust for its neighbor’s confounders in addition to its own.
A3Piecewise smooth invertibility.
The MTP $d(a,l;\delta)$ has an inverse defined
$d^{-1}(a,l;\delta)=\sum_{k=1}^{K(l)}h_{k}(a,l;\delta)\mathbb{I}(a\in\mathcal{A}_{k}(l))\ ,$ (6)
for some countable set of functions $h_{1},\ldots,h_{K(l)}$, where each $h_{k}$ is continuously differentiable with respect to $a$ over the partition $\mathcal{A}_{k}(l)$, a set which may depend on the covariates $l$.
Assumption A3 was first introduced by Haneuse & Rotnitzky (2013) and is standard in the MTP literature (Díaz & van der Laan, 2018; Hejazi et al., 2022; Díaz et al., 2023), where it has been used to ensure the existence of the efficient influence function, allowing for the construction of regular asymptotically linear estimators of the corresponding statistical estimand.
A4Summary coarea.
The coarea of $s_{F_{i}}(\mathbf{a},\mathbf{l})$ exists and exceeds zero almost everywhere. That is, for almost all $\mathbf{a}$, $s_{F_{i}}$ has a Jacobian $J_{\mathbf{a}}$ with respect to $\mathbf{a}$ satisfying
$\sqrt{\det J_{\mathbf{a}}s_{F_{i}}(\mathbf{a},\mathbf{l})J_{\mathbf{a}}s_{F_{i}}(\mathbf{a},\mathbf{l})^{\top}}>0\ .$ (7)
In Equation 7, the left-hand side is called the coarea of $s$. See Negro (2022) for a discussion of coarea in statistics. Together, Assumptions A3 and 7 ensure that the causal estimand can be expressed as an estimable function of $A_{i}^{s}$ and $L_{i}^{s}$ instead of the entire vectors $\mathbf{A}$ and $\mathbf{L}$, and that its efficient influence function exists, which permits the construction of regular asymptotically linear estimators capable of achieving the semi-parametric efficiency bound.
Under Assumptions A17, the counterfactual mean of an induced MTP $\Psi_{n}(\mathsf{P})$ is identified by
$\psi_{n}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\mathsf{P}}(m(A_{i}^{s},L_{i}^{s})\cdot r(A_{i}^{s},A_{i}^{s\circ d},L_{i}^{s})\cdot w(\mathbf{A},\mathbf{L},i))\ ,$ (8)
where, letting $J_{\mathbf{a}}$ denote the Jacobian with respect to $\mathbf{a}$, the nuisances $m$, $r$, and $w$ are defined:
$m(a^{s},l^{s})=\mathbb{E}_{Y}(Y\mid A_{i}^{s}=a^{s},L_{i}^{s}=l^{s})$ (Conditional Mean)
$r(a^{s},a^{s\circ d^{-1}},l^{s})=\frac{p(a^{s\circ d^{-1}}\mid l^{s})}{p(a^{s}\mid l^{s})}$ (Density Ratio)
$w(\mathbf{a},\mathbf{l},i)=\sqrt{\frac{\det J_{\mathbf{a}}(s_{F_{i}}\circ d^{-1})(\mathbf{a},\mathbf{l};\delta)J_{\mathbf{a}}(s_{F_{i}}\circ d^{-1})(\mathbf{a},\mathbf{l};\delta)^{\top}}{\det J_{\mathbf{a}}s_{F_{i}}(\mathbf{a},\mathbf{l})J_{\mathbf{a}}s_{F_{i}}(\mathbf{a},\mathbf{l})^{\top}}}$ (Induced MTP Weights)
See Supplementary Material LABEL:supp:section:identification for a proof. While $m$ and $r$ are nuisance parameters that must be estimated from the data, $w$ is a deterministic function of the data and the choices of $s$ and $d$ made by the investigator. Since $m$ and $r$ only depend on the summaries of $\mathbf{A}$ and $\mathbf{L}$, this identification result effectively factorizes the estimand: all estimation can proceed using exposure and confounder summaries instead of their individual-level values. Furthermore, while the form of $w$ may appear complex, if $s$ and $d$ are linear, $w$ often simplifies considerably. For example, if $d(a,l;\delta)=\delta\cdot a$ and $s_{F_{i}}(\mathbf{a},\mathbf{l})=\sum_{j\in F_{i}}\omega_{j}a_{j}$ (a weighted sum), then $w(\mathbf{a},\mathbf{l},i)=1/\delta$. Unlike for the estimand of Ogburn et al. (2022), no Monte Carlo procedures are necessary to estimate these nuisance quantities.
3.3 Asymptotically Efficient Estimation
It is well-established that standard plug-in and re-weighting estimators cannot leverage flexible machine learning or non-parametric regression strategies for estimation of nuisance parameters without incurring possibly severe asymptotic bias (Koshevnik & Levit, 1977; Pfanzagl & Wefelmeyer, 1985; Bickel et al., 1993). Constructing a consistent estimator that achieves the semi-parametric efficiency bound—the lowest possible variance among regular asymptotically linear estimators—in the non-parametric model $\mathcal{M}$ requires alternative strategies, for which the efficient influence function is a common ingredient. Such strategies include one-step bias-corrected estimation (Pfanzagl & Wefelmeyer, 1985; Bickel et al., 1993), unbiased estimating equations (van der Laan & Robins, 2003), targeted maximum likelihood (or minimum loss) estimation (van der Laan & Rubin, 2006; van der Laan & Rose, 2011), and, most recently, double machine learning (Chernozhukov et al., 2018). In many instances, these distinct frameworks yield doubly robust estimators.
We use doubly robust to mean that, for nuisances $m$ and $r$, consistent estimation occurs if the nuisance error product $\lVert\hat{m}-m\rVert_{2}\lVert\hat{r}-r\rVert_{2}$ is asymptotically negligible. Many common non-parametric regression algorithms can be shown to converge at $n^{1/4}$ rates under certain assumptions (see, e.g., Section 4.3 of Kennedy (2024) or Section 4.1 of Díaz et al. (2021) and references therein). Consequently, when such flexible algorithms are used for nuisance estimation, the product of their convergence rates will be at least $n^{1/2}$ (or faster), making a key second-order remainder term in the von Mises expansion of the estimator and target parameter negligible; moreover, this allows for a doubly robust estimator to achieve consistency under misspecification of either of two nuisance estimators. For a doubly robust estimator to achieve the semi-parametric efficiency bound, both nuisance estimators must be correctly specified (see, e.g., van der Laan & Rose, 2011; Kennedy, 2024).
To construct semi-parametric efficient estimators of $\psi_{n}$, we build upon theoretical developments for estimating the causal effect of a stochastic intervention under interference. Assuming the SCM (4), Sofrygin & van der Laan (2017) showed that such a causal effect may be identified as
$\psi_{n}^{\star}=\frac{1}{n}\sum_{i=1}^{n}\int_{\mathcal{L}^{s}}\int_{\mathcal{A}^{s}}m(a_{i}^{s},l_{i}^{s})\bar{p}^{\star}(a_{i}^{s}\mid l_{i}^{s})p(l_{i}^{s})\partial\mu(a_{i}^{s},l_{i}^{s})\ ,$ (9)
where $\bar{p}^{\star}$ is the mixture of neighbors’ conditional exposure distributions that results from replacing each $A_{i}$ with a random draw $A_{i}^{\star}$ from some user-specified distribution. As noted by other authors (Young et al., 2014; Díaz & van der Laan, 2018), an MTP may be expressed as a variant of a stochastic intervention whose replacement density depends on $A^{s}$ and satisfies piecewise smooth invertibility. Consequently, an induced MTP is a stochastic intervention where
$\bar{p}^{\star}(a_{i}^{s}\mid l_{i}^{s})=p_{A^{s}}(a_{i}^{s\circ d^{-1}}\mid l_{i}^{s})\sqrt{J_{\mathbf{a}}(s_{F_{i}}\circ d^{-1})(\mathbf{a},\mathbf{l};\delta)J_{\mathbf{a}}(s_{F_{i}}\circ d^{-1})(\mathbf{a},\mathbf{l};\delta)^{\top}}\ ,$ (10)
which follows from the change-of-variables formula for functions whose Jacobians $J_{\mathbf{a}}$ are not square. We show this equivalence in Lemma LABEL:supp:lemma:change-of-variables and Section LABEL:supp:section:eif of the Supplementary Material.
The efficient influence function (EIF) for regular asymptotically linear estimators of the statistical functional that identifies the causal effect of a stochastic intervention under interference was first derived by van der Laan (2014). Treating the EIF as an estimating equation is a standard strategy for the construction of semi-parametric efficient estimators (Pfanzagl & Wefelmeyer, 1985; Bickel et al., 1993; van der Laan & Rose, 2011). Applying the representation in Equation (10), we prove in Section LABEL:supp:section:eif of the Supplementary Material that the EIF for the causal effect of an induced MTP is
$\bar{\phi}_{\mathsf{P}}=\frac{1}{n}\sum_{i=1}^{n}r(A_{i}^{s},A_{i}^{s\circ d^{-1}},L_{i}^{s})w(\mathbf{A},\mathbf{L},i)(Y_{i}-m(A_{i}^{s},L_{i}^{s}))+\mathbb{E}_{\mathsf{P}}[m(A_{i}^{s\circ d},L_{i}^{s})\mid\mathbf{L}=\mathbf{l}]-\psi_{n}\ .$ (11)
The EIF $\bar{\phi}_{\mathsf{P}}$ is expressed as an empirical mean because it represents the influence of all $n$ units within the single draw of the interference network that is observed. Despite only observing a single realization of the network, it is clear that this EIF is composed of individual $i$-specific components, with the corresponding EIF estimating equation admitting the expression:
$\bar{\phi}_{\mathsf{P}}=\frac{1}{n}\sum_{i=1}^{n}\phi_{\mathsf{P}}(O_{i})-\psi_{n}=0\ ,$ (12)
where $\phi_{\mathsf{P}}(O_{i})=r(A_{i}^{s},A_{i}^{s\circ d^{-1}},L_{i}^{s})w(\mathbf{A},\mathbf{L},i)(Y_{i}-m(A_{i}^{s},L_{i}^{s}))+\mathbb{E}_{\mathsf{P}}[m(A_{i}^{s\circ d},L_{i}^{s})\mid\mathbf{L}=\mathbf{l}]$. Ogburn et al. (2022) exploit this structure to prove the following theorem, reproduced for convenience:
Theorem 1 (Central Limit Theorem of Ogburn et al. (2022)).
Suppose an estimator $\hat{\psi}_{n}$ is a solution to the EIF estimating equation of Equation (12) and that $K_{\text{max}}^{2}/n\rightarrow 0$ as $n\rightarrow\infty$, where $K_{\text{max}}$ denotes the maximum node degree in the network. Under the SCM (4) and mild regularity conditions, including that the estimating function is bounded,
$\sqrt{C_{n}}(\hat{\psi}_{n}-\psi_{n}^{\star})\overset{d}{\rightarrow}N(0,\sigma^{2})\ ,$ (13)
for some finite $\sigma^{2}$ and some constant $C_{n}$ such that $n/K_{\text{max}}^{2}\leq C_{n}\leq n$ .
This theorem states that if $K_{\text{max}}$ grows asymptotically no faster than $n^{1/2}$, then an estimator $\hat{\psi}_{n}$ of $\psi_{n}^{\star}$ attains a normal limiting distribution centered about $\psi_{n}^{\star}$. In addition, these authors argue, based on earlier results of van der Laan (2014), that such an estimator of $\psi_{n}^{\star}$ will exhibit double robustness with respect to the nuisance estimators $\hat{m}$ and $\hat{r}$, with a second-order remainder term of the form $\lVert\hat{m}-m\rVert_{2}\lVert\hat{r}-r\rVert_{2}=o_{\mathsf{P}}(\sqrt{C_{n}})$, where the estimator $\hat{\psi}_{n}$ of $\psi_{n}^{\star}$ remains consistent as long as the rate-product of the differences of the nuisance estimators $\{\hat{m},\hat{r}\}$ from their respective targets $\{m,r\}$ converges at the specified rate. Since our target estimand $\psi_{n}$ is equivalent to $\psi^{\star}_{n}$ under an MTP and appropriate identification conditions, using Theorem 1 and the EIF in Equation (11), we can construct semi-parametric efficient estimators of $\psi_{n}$ in at least two ways, which we outline next.
One-Step Estimation. A one-step bias-corrected estimator (Pfanzagl & Wefelmeyer, 1985; Bickel et al., 1993) uses Equation (12) to de-bias an initial plug-in estimator by adding to it an estimated $\phi_{\hat{\mathsf{P}}_{n}}(O_{i})$, constructed based on nuisance estimators. While we use $\mathsf{P}_{n}$ to denote the empirical distribution, we use $\hat{\mathsf{P}}_{n}$ to denote the empirical distribution augmented by relevant nuisance estimators. The outer expectation in the term $\mathbb{E}_{\mathsf{P}}[\hat{m}(A_{i}^{s\circ d},L_{i}^{s})\mid\mathbf{L}=\mathbf{l}]$ in the EIF can be dropped because the sample means of $\hat{m}(A_{i}^{s\circ d},L_{i}^{s})$ and $\mathbb{E}_{\mathsf{P}}[\hat{m}(A_{i}^{s\circ d},L_{i}^{s})\mid\mathbf{L}=\mathbf{l}]$ both converge to the same value (see Section LABEL:supp:section:one-step of the Supplementary Material for a proof). Then, the one-step (OS) estimator is
$\hat{\psi}_{n}^{\text{OS}}=\frac{1}{n}\sum_{i=1}^{n}\phi_{\hat{\mathsf{P}}_{n}}(O_{i})\ ,$ (14)
where $\phi_{\hat{\mathsf{P}}_{n}}(O_{i})=\hat{r}(A_{i}^{s},A_{i}^{s\circ d^{-1}},L_{i}^{s})w(\mathbf{A},\mathbf{L},i)(Y_{i}-\hat{m}(A_{i}^{s},L_{i}^{s}))+\hat{m}(A_{i}^{s\circ d},L_{i}^{s})$. Since the estimating functions are no longer centered, variance estimates must be adjusted (as described in Section 3.5).
Targeted Maximum Likelihood Estimation (TMLE). Although the one-step estimator is semi-parametric efficient, it is not a substitution estimator: it may yield estimates outside the bounds of the parameter space. TMLE—a general template for the construction of substitution estimators that appropriately solve the EIF estimating equation (van der Laan & Rubin, 2006; van der Laan & Rose, 2011)—resolves this shortcoming. Estimation under an induced MTP proceeds as follows:
  1. 1.
    Estimate nuisances functions to obtain $\hat{r}(A_{i}^{s},L_{i}^{s})$, $\hat{r}(A_{i}^{s\circ d},L_{i}^{s})$ $\hat{m}(A_{i}^{s},L_{i}^{s})$, and $\hat{m}(A_{i}^{s\circ d},L_{i}^{s})$, and compute weights $w(\mathbf{A},\mathbf{L},i)$ and $w(\mathbf{A}^{d},\mathbf{L},i)$ from the data based on the chosen $s$ and $d$.
  2. 2.
    Fit a one-dimensional parametric fluctuation model regressing $\hat{r}(A_{i}^{s},L_{i}^{s})w(\mathbf{A},\mathbf{L},i)$ on $Y_{i}$ with offset $\hat{m}(A_{i}^{s},L_{i}^{s})$; a common choice for this is logistic regression $\text{logit}(Y_{i})=\text{logit}(\hat{m}(A_{i}^{s},L_{i}^{s}))+\varepsilon\hat{r}(A_{i}^{s},L_{i}^{s})w(\mathbf{A},\mathbf{L},i)$, where the $Y_{i}$ are rescaled to ensure each lies in the open unit interval $(0,1)$.
  3. 3.
    Compute $\hat{m}^{\star}(A_{i}^{s\circ d},L_{i}^{s})=\text{expit}(\text{logit}(\hat{m}(A_{i}^{s\circ d},L_{i}^{s})+\hat{\varepsilon}\hat{r}(A_{i}^{s\circ d},L_{i}^{s})w(\mathbf{A}^{d},\mathbf{L},i)$ if using the logistic fluctuation model, or more broadly, compute the corresponding parametric fluctuation model predictions under the MTP intervention, based on the MLE $\hat{\varepsilon}$ of $\varepsilon$.
  4. 4.
    Compute the TML estimate as $\psi_{n}^{\text{TMLE}}=\frac{1}{n}\sum_{i=1}^{n}\hat{m}^{\star}(A_{i}^{s\circ d},L_{i}^{s})$.
In Step 2, the corresponding estimate $\hat{\varepsilon}$ is used to fluctuate the initial estimator $\hat{m}(A_{i}^{s},L_{i}^{s})$, using the summary-weighted density ratio $\hat{r}(A_{i}^{s},L_{i}^{s})w(\mathbf{A},\mathbf{L},i)$, to an updated version $\hat{m}^{\star}(A_{i}^{s},L_{i}^{s})$ in such a way that the EIF estimating equation is solved. Logistic regression is commonly employed to guarantee each prediction $\hat{m}^{\star}$ falls in the bounds of the parameter space, thus respecting global constraints (Gruber & van der Laan, 2010). One can also fit an intercept-only model with $\hat{r}(A_{i}^{s},L_{i}^{s})w(\mathbf{A},\mathbf{L},i)$ as weights themselves, which may improve stability. In what follows, we refer to this estimator as “network-TMLE,” following nomenclature introduced by Zivich et al. (2022), who discuss its use in practical settings and examine its properties in simulation experiments.
3.4 Nuisance Parameter Estimation
The underlying form of nuisance parameters $m$ and $r$ is usually recognized as being unknown and may involve, for example, complex nonlinear interactions between covariates. Therefore, it is desirable to estimate such quantities using flexible regression or machine learning approaches. This process often employs $K$-fold cross-fitting (Zheng & van der Laan, 2010; Chernozhukov et al., 2018) to eliminate asymptotic bias arising from empirical process terms (Pfanzagl & Wefelmeyer, 1985; Klaassen, 1987).
Perhaps surprisingly, the form of cross-fitting used for independent units still accomplishes the same goals even in our network setting. That is, splitting the data into $K$ folds, fitting nuisance functions on each combination of K1K-1 folds, and obtaining out-of-sample predictions for each held-out fold is sufficient to render any empirical process bias asymptotically negligible, even when units may be correlated according to Equation 4. A formal proof can be found in Section LABEL:supp:section:crossfitting of the Supplementary Material. At a high level, it follows similarly to an $m$-dependence argument: a sample mean is unbiased even under dependence, and its variance will shrink at a rate dependent on the maximum degree of the network, which in our setting is limited by assumption. The result mirrors similar numerical findings in the longitudinal data setting (e.g., Fuhr & Papies, 2024). Consequently, cross-fitting can be performed identically whether units may be independent or correlated according to a network whose maximum degree grows according to Theorem 1.
Since the outcome regression $m$ is a conditional expectation function, it can be estimated using any supervised learning algorithm. We recommend super learning (van der Laan et al., 2007), that is, fitting an ensemble of candidate regression algorithms in a cross-validated manner and selecting the candidate yielding the lowest cross-validated risk (for continuous outcomes, we use mean-squared error). This guarantees that the selected algorithm performs asymptotically as well as the best candidate algorithm in the library used to construct the ensemble (van der Laan et al., 2004; van der Vaart et al., 2006), even in the fixed regression design setting presently considered (Davies & van der Laan, 2016). Phillips et al. (2023) proposed guidance and heuristics to aid in overcoming the considerable challenge of assembling a candidate library from the diversity of learning algorithms available.
An advantage of the MTP framework is that the implied form of the intervention, unlike general stochastic interventions (Díaz & van der Laan, 2012; Sofrygin & van der Laan, 2017; Ogburn et al., 2022), facilitates direct estimation of the density ratio $r$, circumventing the need to learn a conditional density function. A common density ratio estimator is based on probabilistic classification (Qin, 1998; Cheng & Chu, 2004), in which a classifier is trained to distinguish between natural and intervened samples, and its output is transformed into a density ratio using Bayes’ rule. The use of this method with an MTP is described in detail by Díaz et al. (2023, see Section 5.4). We recommend super learning with binary log loss to select the optimal classifier. One can also employ kernel-based methods, including kernel mean matching, Kullback-Leibler importance estimation, and least-squares importance fitting, as outlined in detail by Sugiyama et al. (2012).
3.5 Variance Estimation
Summary measures are correlated if they aggregate the same neighboring units. As a result, the as-iid sample variance of the estimated EIF will be conservative, especially if the network degree distribution is highly skewed (Sofrygin & van der Laan, 2017). A consistent variance estimator for the solution to an arbitrary estimating equation $\frac{1}{n}\sum_{i=1}^{n}\varphi_{i}=0$ with a centered estimating function $\varphi_{i}$ satisfying $\mathbb{E}[\varphi_{i}]=0$ is given by
$\hat{\sigma}^{2}=\frac{1}{n^{2}}\sum_{i,j}G(i,j)\varphi_{i}\varphi_{j}\ ,$ (15)
where $G(i,j)=1$ if $i=j$, $i$ and $j$ are in each others’ dependency neighborhoods, or $i$ and $j$ share a friend $k$, and $G(i,j)=0$ otherwise. A useful feature of $\hat{\sigma}^{2}$ is that it automatically incorporates the scaling factor $C_{n}$ based on the CLT of Ogburn et al. (2022). Intuitively, this is because $\hat{\sigma}^{2}$ includes $n$ “variance” terms $\varphi_{i}^{2}$, as well as a certain number of “covariance” terms $\varphi_{i}\varphi_{j}$ (for $i\neq j$), the number of which scales with $n/K_{\max}^{2}\leq C_{n}\leq n$, representing the rate of connectivity in the network. Consequently, $(1-\alpha)$% Wald-style confidence intervals can be constructed via the standard approach: $\hat{\psi}_{n}\pm\Phi^{-1}(\alpha/2)\sqrt{\hat{\sigma}^{2}/n}$, where $\Phi$ is the CDF of the standard normal distribution.
[…]
$\hat{\psi}_{n}(\lvert F_{i}\rvert)=\frac{1}{|\mathcal{N}(|F_{i}|)|}\sum_{j\in\mathcal{N}(\lvert F_{i}\rvert)}\phi_{\hat{\mathsf{P}}_{n}}(O_{j})\ ,$ (16)
where $\mathcal{N}(m)=\{k\mathrel{\mathop{\ordinarycolon}}k\in 1,\ldots,n,|F_{k}|=m\}$. Then, letting $\varphi_{i}=\phi_{\hat{\mathsf{P}}_{n}}(O_{i})-\hat{\psi}_{n}(|F_{i}|)$, the estimator $\hat{\sigma}^{2}$ in Equation (15) consistently estimates the variances of $\hat{\sigma}^{\text{OS}}_{n}$ and $\hat{\sigma}^{\text{TMLE}}_{n}$, provided both nuisance estimators are consistent for their targets (see Supplementary Material, Section LABEL:supp:section:variance-estimator, for a proof).
4 Results: Numerical Experiments
We now empirically evaluate the estimators described in Section 3. All data in this section are simulated in the numerical computing language Julia (Bezanson et al., 2017), using the package CausalTables.jl (Balkus & Hejazi, 2025). MTP estimates are computed using ModifiedTreatment.jl (Balkus & Hejazi, 2024), a Julia package implementing the estimators described in this work. Unless otherwise specified, conditional means and density ratios are estimated using super learning—from an ensemble consisting of a GLM, random forest, and multiple gradient boosted tree models with various sets of hyperparameters—to select the algorithm minimizing the cross-validated empirical risk with respect to an appropriate loss (MSE for the conditional mean, binary log-loss for the density ratio). Code for these experiments and the data analysis is available on GitHub at https://github.com/salbalkus/pub-code-mtp-interfere.
4.1 Results of Experiments with Synthetic Data
First, we evaluate estimators on simulated data with both network interference and nonlinear relationships between confounders, exposure, and outcome. We simulate this data using three common network structures: Erdős-Rényi (with $p=3/n$), static scale-free (with $\lambda=3.5)$, and Watts-Strogatz (with $K=6$ and $\beta=0.5$). Data are generated according to the following set of structural equations:
$\begin{aligned} \mathbf{L}_{1}&\sim\text{Beta}(3,2);\mathbf{L}_{2}\sim\text{Poisson}(100);\mathbf{L}_{3}\sim\text{Gamma}(2,4);\mathbf{L}_{4}\sim\text{Bernoulli}(0.6)\\ m_{L}&=\Big(1+L_{4}\Big)\cdot\Big(-2(\mathbbm{I}(L_{1}>0.3)+\mathbbm{I}(L_{2}>90)+\mathbbm{I}(L_{3}>5))-(\mathbbm{I}(L_{1}>0.5)+\\ &\mathbbm{I}(L_{2}>100)+\mathbbm{I}(L_{3}>10))+2(\mathbbm{I}(L_{1}>0.7)+\mathbbm{I}(L_{2}>110)+\mathbbm{I}(L_{3}>15))\Big)\\ \mathbf{A}&\sim\text{Normal}(m_{L}-5,1.0)\,\,\text{and}\,\,\mathbf{A}^{s}=\Big[\sum_{j\in F_{i}}A_{i}\Big]_{i=1}^{n}\\ m_{A}&=-2\mathbbm{I}(A>-2)-\mathbbm{I}(A>1)+3\mathbbm{I}(A>3);m_{A_{s}}=3\mathbbm{I}(A_{s}>0)+\mathbbm{I}(A_{s}>6)+\mathbbm{I}(A_{s}>12)\\ \mathbf{Y}&\sim\text{TruncNormal}(m_{L}\cdot(1+0.2m_{A}+m_{A_{s}})+5,2.0)\ ,\end{aligned}$
where TruncNormal denotes a normal distribution truncated at six standard deviations. We estimate the counterfactual mean under the MTP $d(a,l)=a+0.25$ using network-TMLE. Being very similar, one-step estimator results are omitted for brevity.
Figure 2 demonstrates how the empirical performance of our estimator using network-TMLE aligns with theoretical results outlined in Section 3. Bias reliably converges towards zero with increasing sample size across all three network structures, and is highest for scale-free, the network with the most skewed distribution of node degree. Since $K_{\max}$ in each network grows at roughly $\log(n)$ or slower, we use $\sqrt{C}_{n}=\sqrt{n}/\log(n)$ as a scaling factor for the scaled bias and MSE; these values still decrease, so we can conclude that the convergence rate of network-TMLE matches the expected rate. MSE converges to the efficiency bound; see Supplementary Material LABEL:supp:section:eff-bound-elaboration for details on this bound. The MSE is highest in the Watts-Strogatz simulation, whose network has the highest average node degree (and thus the greatest correlation between units)—hence, a denser network increases variance. Furthermore, the coverage rate of the 95% confidence interval approaches the nominal level for all three graphs. Based on these observations, we conclude that the network-TMLE estimator performs as expected (see Supplementary Material LABEL:supp:section:more-sim-results for a comparison to classical methods).
Figure 2: Asymptotic performance of network-TMLE in simulation on multiple network structures.
4.2 Results of Experiments with Semi-Synthetic Data
In a second simulation experiment, we generate semi-synthetic data based on the real-world dataset that inspired this work and that is analyzed in Section 5. Sixteen normalized confounding variables are taken as fixed covariates, with the 2013 and 2019 LODES commuting pattern data (U.S. Census Bureau, 2024) for California as the network profile $\mathbf{F}$. From these, we draw exposure and outcome from normal distributions with means following a linear function of the confounders and $A^{s}$, a trimmed sum of neighbors’ values of $A$ over neighbors that contributed at least  2.5% of commuters into the given unit. These are represented by structural equations, conditional on $\mathbf{L}$ and $\mathbf{F}$, expressed as
$\begin{aligned} \mathbf{L}_{k}^{s}&=\Big[\sum_{j\in F_{i}}L_{k,i}\Big]_{i=1}^{n}\,\,\mathbf{A}\sim\text{Normal}\Big(\sum_{k=1}^{16}\mathbf{L}_{k}-50,1.0\Big)\,\,\text{and}\,\,\mathbf{A}^{s}=\Big[\sum_{j\in F_{i}}A_{i}\Big]_{i=1}^{n}\\ \mathbf{Y}&\sim\text{TruncNormal}\Big(\mathbf{A}+\mathbf{A}^{s}+\sum_{k=1}^{16}\mathbf{L}_{k}+\sum_{k=1}^{16}\mathbf{L}_{k}^{s}-50,1.0\Big)\ ,\end{aligned}$ (17)
We estimate the population causal effect of an MTP in 400 simulations using four different estimators: network-TMLE with correctly-specified GLMs (linear regression for the outcome regression and logistic regression for the density ratio probabilistic classifier), network-TMLE with super learning for both nuisance estimators, IID-TMLE with correctly-specified GLMs, and main-terms linear regression. In this simulation, $n=1652$, the same as in the dataset considered in Section 5.
Table 1 depicts the operating characteristics of the estimators. It is clear from inspection that accounting for network interference can yield dramatically lower bias, even when the outcome regression is otherwise correctly specified. In this case, using network-TMLE with a correctly specified GLM decreased the mean percent bias from over 20% to about 0.1%, and cut variance by about 25%. Even if super learning is used, the mean percent bias of the network-TMLE remains relatively low, at around 1%, and the variance is approximately identical. Furthermore, confidence interval coverage for the network-TMLE is close to nominal under both configurations of nuisance estimators. Alternative methods suffer greatly reduced coverage due to their severe bias.
Table 1: Network-TMLE performance on semi-synthetic data versus competing estimators
Method Learner Bias (%) Variance Coverage (%) CI Width
Network TMLE Correct GLM $0.11$ $1.56$ $96.2$ $4.88$
Super Learner $1.03$ $1.56$ $94.0$ $4.88$
Classical TMLE Correct GLM $20.43$ $2.11$ $54.8$ $5.70$
Linear Regression $20.62$ $2.12$ $55.0$ $5.71$
5 Illustrative Data Analysis: Mobile Source Air Pollution
In this section, we use induced MTPs to analyze the causal effect of “zero-emissions” vehicle (ZEV) uptake in California on NO2 air pollution using observational data. NO2 is a byproduct of gasoline-powered vehicles shown to be associated with adverse health outcomes, including respiratory issues and mortality (Hesterberg et al., 2009; Gillespie-Bennett et al., 2010; Faustini et al., 2014). In the U.S., national ambient air quality standards (NAAQs) are set by the Environmental Protection Agency (EPA) to limit the amount of allowable NO2 in the air (U.S. Environmental Protection Agency, 2018).
5.1 Applied Science Background and Motivations
As ZEVs do not produce tailpipe emissions, governments have heavily promoted their adoption to limit air pollution. State governments must ascertain whether these policies work as intended; hence, understanding how much localized air pollution has decreased as a result of increased ZEV adoption is of both scientific and policy interest. Garcia et al. (2023) analyzed whether ZEV uptake was associated with NO2 air pollution across ZIP code tabulation areas (ZCTA) in California from 2013 to 2019. They found that an increase of 20 ZEVs per 1000 population units was associated with a reduction of 0.41 parts per billion in NO2 within a given ZCTA on average; however, with a $p$-value of 0.252 and confidence interval of (1.12,0.29)(-1.12,0.29), this finding was far from statistically significant.
This is a problem in which network interference arises as a result of people driving outside the ZCTA in which they reside; that is, individuals commute in their vehicles to other ZCTAs, emitting pollution there as well. Garcia et al. (2023) controlled for confounding using linear regression, without accounting for interference due to individuals commuting in vehicles from one ZCTA unit to another. As evidenced by our simulation experiments in Section 4, neglecting interference can induce bias and yield misleading confidence intervals. In addition, overreliance on restrictive parametric modeling may be cause for concern, as linearity imposes a strict functional form which may not adequately capture the underlying complexity, resulting in model misspecification bias. The goal of our data analysis is to study the same research question as Garcia et al. (2023) and compare how flexibly evaluating the causal effect of an induced MTP can improve inference in this setting.
5.2 Estimation Strategy
Following Garcia et al. (2023), we compute the exposure, the percentage of light-duty vehicles in California actively registered as ZEVs by April 2019, for each ZCTA from the California Energy Commission (2024). The outcome, change in NO2 from 2013 to 2019 in parts per billion (ppb), is spatially aggregated from a 1 km grid of estimated pollution levels (Cooper, 2022a, b); this provides finer-grain ZCTA-level estimates than the raw sensor data used by Garcia et al. (2023). Socioeconomic confounders are accessed from the U.S. Census via Walker & Herman (2024), while land-use confounders are obtained from the U.S. Environmental Protection Agency (2024); see Section LABEL:supp:section:data-sources of the Supplementary Material for summary statistics. We performed spatial alignment and areal weighted interpolation of missing values using the sf (Pebesma, 2018) and areal (Prener et al., 2019) packages in the R language for statistical computing and graphics (R Core Team, 2025).
Finally, to account for interference, we rely on commuting networks describing the number of people traveling between home and work ZCTAs (de Souza et al., 2023; U.S. Census Bureau, 2024). The induced MTP summary estimates the percentage of incoming work commuters driving ZEVs during the 2013–2019 time period, as used for the outcome. This was computed by summing the percentage of ZEVs in each neighboring ZCTA in each year’s network, with each unit’s summand weighted by the percentage of its neighbor’s population who drive to work and normalized. Sums of neighboring covariates are also included as potential confounders for the induced MTP.
We compared the following data-analytic strategies: (1) estimating an induced MTP using network-TMLE, (2) estimating a classical MTP using IID-TMLE, and (3) estimating a main-terms GLM regression coefficient. Each seeks to answer the same question: “how much more would average NO2 have decreased from 2013 to 2019 had each ZCTA experienced an increase in the percentage of vehicles registered as ZEV by 2019?” Strategy (1) accounts for interference and estimates nuisances using flexible super learning. Strategy (2) uses super learning but ignores interference. Strategy (3) imposes a strict linearity assumption and ignores interference. Both super learning procedures selected random forest for the outcome regression and density ratio probabilistic classifier.
5.3 Analytic Results
Figure 3 displays the three estimates of additive and multiplicative effects of an MTP increasing the proportion of ZEVs. According to the induced MTP effect estimate, adding 1% to the proportion of ZEVs across ZCTAs would be expected to yield a 0.044 ppb decrease in NO2 on average. This estimate is over 1.3 times larger than the MTP effect estimate arrived at when not accounting for interference (0.032 ppb), and about 2.8 times larger than the GLM-based estimate (0.015 ppb). Similarly, the induced MTP effect estimate indicated that scaling the proportion of ZEVs by 20% across ZCTAs would be expected to yield a 0.048 ppb decrease in NO2 on average—over 1.4 times larger than the naive MTP effect estimate (0.033 ppb) and the GLM-based estimate (0.027 ppb). Interpreted in a scientific context, for the additive shift, the estimated effect based on the induced MTP indicates that ZEV uptake would have accounted for about 7% of the overall mean decrease in NO2 from 2013-2019, whereas a GLM-based estimate would have accounted for only 2.5%.
Figure 3: Estimated effect sizes measuring the expected difference in NO2 across California ZCTAs caused by two different increases in the proportion of ZEVs in 2019.
Figure 4: Estimated effect sizes measuring the expected difference in NO2 across California ZCTAs caused by various additive or multiplicative shifts in the percentage of ZEVs in 2019. Confidence bands use a conservative Bonferroni multiplicity correction.
Accounting for interference and using flexible regression yielded larger effect estimates than those recovered by classical analyses, suggesting commuting contributes significantly to vehicular NO2 pollution. This was especially pronounced for the multiplicative shift, indicating that ZCTAs with many out-commuters exacerbate NO2. Furthermore, while one might expect that accounting for interference would yield wider confidence intervals due to correlation between units, we observe the opposite—drastically lower variance. We conjecture that improvements in estimation of the outcome regression drive this result, as including exposure summaries as covariates reduces model misspecification, allowing the model to explain a larger proportion of the outcome variance.
Figure 4 displays effect estimates over a grid of possible additive and multiplicative shifts. At low-magnitude shifts, the estimates were roughly the same. However, at larger shifts—about 0.75% on the additive scale and 10% on the multiplicative scale—the estimated effects diverged, with those accounting for interference becoming more pronounced. The trend was also less erratic under the multiplicative shift, possibly because it only produces a large shift among ZCTAs with a large proportion of ZEVs, thereby avoiding destabilizing empirical positivity issues. Such a result highlights the importance of considering MTPs that can be easily estimated from the data available.
Our analysis focused on NO2 since it is a well-understood pollutant. That said, from 2013–2019, average NO2 levels changed little across ZCTAs, as all ZCTAs within California have remained well within limits regarded as safe based on EPA guidance (U.S. Environmental Protection Agency, 2018), so the effects are inherently small. However, suppose policymakers sought further reductions and would impose a ZEV-promoting policy given statistically significant evidence. Then, according to Figures 3 and 4, evidence from an induced MTP would have resulted in imposing a ZEV-promoting policy ($p<0.05$) versus not imposing such a policy ($p>0.05$) if interference was ignored. Hence, our method provides much stronger and more conclusive evidence regarding even the small effect sizes in this setting.
6 Discussion
In this work, we introduced the induced modified treatment policy, a new class of MTP that accounts for known network interference. This intervention is useful for causal inference in observational data settings that feature continuous exposures and network interference. We established identification of the causal effect of an induced MTP using a novel application of the coarea formula and outlined procedures for constructing semi-parametric-efficient estimators capable of incorporating flexible nuisance estimation strategies via, for example, machine learning or non-parametric regression.
Using simulation experiments, we showed that interference can result in significant bias when it is not corrected using an induced MTP. Our illustrative data analysis demonstrates the perils of ignoring interference and applying restrictive parametric modeling strategies with observational spatial data, as is often done in environmental epidemiology and related fields. Such oversimplifications can suggest starkly different scientific conclusions than those provided by our proposed strategy.
In practice, several challenges remain. One limitation is that ratios of conditional densities are still more difficult to estimate than conditional expectation functions, especially in high-dimensional settings (Sugiyama et al., 2012). This problem is exacerbated under interference: for units with an excessively large number of neighbors, the density ratio between the natural and post-intervention exposure could grow extremely large, destabilizing downstream estimates. More modern “balancing” tools, such as Riesz regression (Chernozhukov et al., 2021), may help overcome this issue.
In addition, practitioners must know how the interference arises. This does often occur: in our data analysis, the form of interference arose naturally as a part of the scientific question that considered an intervention on all vehicles entering a given ZCTA, not just those registered there. If an investigator has reason to suspect interference, they may also suspect the underlying process by which it occurs. However, there are cases where an investigator may not know the form. Although discussed in recent work (Hoshino & Yanagi, 2023; Ohnishi et al., 2022), this remains an avenue for future research.
Other potential areas of future research involve extending the induced MTP to more complex settings. In time-varying data, for example, one might consider using an induced longitudinal MTP (Díaz et al., 2023), which would involve sequential regression-based algorithms to account for summary measures of exposures subject to time-varying confounding in a network profile that may itself evolve across time. Further work overcoming the technical limitations of estimating the effects of induced MTPs will be critical to facilitate answering more complex causal inference questions in settings where continuous exposures are measured in datasets exhibiting network interference.
Acknowledgments
SVB was supported in part by grants from the National Institute of Environmental Health Sciences (award no. T32 ES007142) and the National Science Foundation (award no. DGE 2140743). The authors thank Rachel Nethery for her help in initiating the motivating data analysis.
Conflicts of interest: The authors have no conflicts of interest to disclose.